OpenStudy (anonymous):

solve the system using elimination { 2x+6y=-12 { 5x-5y=10 A. (2,1) B. (0,-2) C. (-2,0) D (1,2)

OpenStudy (anonymous):

2x+6y=-12--------(1 5x-5y=10--------(2 multiplying (1 by 5 and (2 by 6 and adding them we find 10x+30y=-60--------(1 30x-30y=60---------(2 ---------------------- 40x=0 x=0/40 x=0 Now substituting x=0 in eq (1 we find 2*0+6y=-12 0+6y=-12 6y=-12 y=-12/6 y=-2 {x,y}={0, -2} Thus option B. (0,-2) is correct Ans.

OpenStudy (anonymous):

thank you :) that really helps alot